how much soil containing 45% sand do you need to add to 1 ft3 of soil containing 15% sand in order to make a soil containing 35%?
Answer by stanbon(75887)

how much soil containing 45% sand do you need to add to 1 ft3 of soil containing 15% sand in order to make a soil containing 35%?
-------
Equation:
sand + sand = sand
0.45x + 0.15*1 = 0.35(x+1)
---
Multiply thru by 100:
45x + 15 = 35x + 35
10x = 10
x = 1 ft^3 (amount of 45% soil needed for the mixture)
================
Cheers,
Stan H.
================